OpenStudy (anonymous):

Solve using equation using Zero Product Property. (2x+6)(3x-6)=0 A.3 and 2 B.-3 and -2 C.-3 and 2 D.3 and -2

OpenStudy (mathstudent55):

The zero product property is that if a product equals zero, then one of its factors equals zero. Using the zero product property means that you set each factor equal to zero, and find which values of x cause the factors to equal zero. Solve 2x + 6 = 0 for x, and solve 3x - 6 = 0 for x.

OpenStudy (anonymous):

Then after that?

OpenStudy (mathstudent55):

Can you solve 2x + 6 = 0 for x?

OpenStudy (anonymous):

Honestly I would have to look it up

OpenStudy (mathstudent55):

Ok, let's go over it together. Here is what solving your problem should look like.

OpenStudy (mathstudent55):

\((2x+6)(3x-6)=0\) \(2x + 6 = 0 ~~~~~or~~~~~3x - 6 = 0\) \(2x = -6~~~~~~~~~or~~~~~~~3x = 6\) \(x = -3~~~~~~~~~~~or~~~~~~~~x = 2\)

OpenStudy (anonymous):

That was so easy!

OpenStudy (mathstudent55):

The first step is to separate the product equaling zero into two simpler equation, each one equaling zero. Then we solve each equation. For 2x + 6 = 0, first we subtract 6 from both sides, then we divide both sides by 2. For 3x - 6 = 0, first we add 6 to both sides, then we divide both sides by 3. Then we write teh solutions to both equations separated by the word "or".

OpenStudy (mathstudent55):

Now that you see the method, perhaps next time you'll be able to do it yourself.

OpenStudy (anonymous):

Yeah I get how to do it now. Thank you very much!!

OpenStudy (mathstudent55):

You're welcome.